A 500 line/mm diffraction grating is illuminated by light of wavelength 510nm. how many diffraction orders are seen, and what is the angle of each?
Answer #1

the spacing between the lines d = 1mm/500 = 2000 nm

sinthetam = m*lambda/d = m*510nm/2000nm = 0.255m

thetam = sin^-1(0.255m)

theta0 = sin^-1(0.255*0) = 0

theta1 = 14.77

theta2 = 30.66

theta3 = 49.9

theta4 = does not exist

There are three orders of bright lines on each side of the central bright line.